On an icy road, a 10-kg cart is given a push that imparts to it an initial speed of v0 = 2.0 m/s. The coefficient of kinetic friction between cart and ice is k = 0.10. Find the distance the cart moves before coming to rest.


1
Expert's answer
2021-05-07T11:18:24-0400

The initial kinetic energy of the cart is:


"K_i = \dfrac{mv_0^2}{2}"

where "m = 10kg" is the mass of the cart, and "v_0 = 2m\/s" is its initial speed. The final kinetic energy is "K_f = 0", since the cart stops.

According to the work-energy theorem, the total work done by friction is:


"W = -\Delta K = -(K_f - K_i) = \dfrac{mv_0^2}{2}"

On the other hand, by definition, the work is:


"W = F_{fr}\cdot d"

where "d" is the distance the cart moves before coming to rest, and "F_{fr}" is the friction force. Int turn the friction force is defined as follows:


"F_{fr} = k N = k mg"

where "k = 0.1" is the coefficient of kinetic friction between cart and ice, and "N = mg" is the normal force (equal to the weight of the car, since it does not move verticaly). Finally, obtain:


"k mg\cdot d = \dfrac{mv_0^2}{2}\\
d = \dfrac{mv_0^2}{2kmg} = \dfrac{v_0^2}{2kg} \\
d = \dfrac{2^2}{2\cdot 0.1\cdot 9.8} \approx 2.0m"

Answer. 2 m.